# Abdoul Kader Kaboré

### v3

@@ Abstract
This study presents the first large-scale empirical evaluation of LLM-generated unit tests at the class level, systematically analyzing four state-of-the-art models—GPT-3.5, GPT-4, Mistral 7B, and Mixtral 8x7B—against EvoSuite across 216,300 test cases from Defects4J, SF110, and CMD (a dataset mitigating LLM training data leakage). We evaluate five prompting techniques—Zero-Shot Learning (ZSL), Few-Shot Learning (FSL), Chain-of-Thought (CoT), Tree-of-Thought (ToT), and Guided Tree-of-Thought (GToT)—assessing syntactic correctness, compilability, hallucination-driven failures, readability, code coverage metrics, and fault detection capabilities.
Our findings challenge prior claims that in-context learning is ineffective for test generation in code-specialized LLMs. Reasoning-based prompting—particularly GToT—significantly enhances test reliability, compilability, and structural adherence in general-purpose LLMs. However, hallucination-driven failures remain a persistent challenge, manifesting as non-existent symbol references, incorrect API calls, and fabricated dependencies, resulting in high compilation failure rates (up to 86%). Execution-based classification and mutation testing reveal that many failing tests stem from hallucinated dependencies, limiting effective fault detection.
Keywords:
Automatic Test Generation · Unit Tests · Large Language Models · Prompt Engineering · Empirical Evaluation
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Fan et al., 2023) have emerged as a promising alternative for test generation, potentially addressing the fundamental limitations of traditional approaches. Initial studies suggest LLM-generated tests feature more natural naming conventions, human-like assertions, and better readability than SBST alternatives (Siddiq et al., 2024a; Yuan et al., 2023; Tang et al., 2024). Moreover, LLMs can generate both functional and exploratory tests that go beyond mere structural coverage (Wang et al., 2024)—a flexibility that aligns with comprehensive testing practices.
Despite their promise, LLM-driven test generation remains largely unproven for production use. Current evaluations focus primarily on syntactic correctness rather than practical integration and execution outcomes. Our research identifies and addresses five critical gaps that must be resolved before LLM-based testing can be reliably deployed:
@@
  • Structural Correctness & Extractability of LLM-Generated Tests: Prior work examined compilability (Siddiq et al., 2024a; Tang et al., 2024; Yang et al., 2024) but not structural correctness and extractability. We apply MSR and CSR metrics (Macedo et al., 2024) to assess whether tests are valid JUnit cases and can be seamlessly integrated into testing pipelines.
  • Execution Reliability and Fault Detection: SBST tests maximize coverage but often lack valid assertions (Fraser and Arcuri, 2011; Shamshiri et al., 2015). LLM tests show high failure rates (30.4% for ChatGPT) (Yang et al., 2024), reducing executable tests for bug detection. ChatGPT detects bugs in only 21% of cases versus 26% for EvoSuite (Yang et al., 2024). We address gaps in prior work by categorizing test errors and applying mutation testing with Pitest to assess assertion validity and defect detection.
@@
This paper: We systematically evaluate LLM-generated test suites and prompt engineering impact. Building on prior comparisons (Tang et al., 2024; Siddiq et al., 2024a; Yang et al., 2024), surveys (Wang et al., 2024), and AI risk studies (Sallou et al., 2023), we conduct a large-scale evaluation across four LLMs and five prompting techniques against EvoSuite. Unlike method-level approaches in prior work, our class-level evaluation aligns with real-world testing workflows.
@@
  3. Systematic Execution Outcome & Fault Detection Analysis: We classify test outcomes as Passed, Failed, or Timeout, revealing gaps between syntactic correctness and execution behavior. We apply mutation testing to assess fault detection by introducing synthetic faults and measuring their detection rate. However, most LLM-generated Passed tests proved unsuitable for this evaluation as they primarily targeted interfaces, abstract classes, or trivial methods without engaging actual program logic. This resulted in a 0% mutation score, revealing that while LLMs can generate syntactically correct tests that pass, these tests often lack engagement with core program logic—making them unsuitable for mutation testing and limiting their ability to verify meaningful program behavior.”
@@
The remainder of this paper is structured as follows: Section 2 covers LLM-based test generation, Section 3 outlines our experimental setup, Section 4 presents evaluation results, Section 5 discusses threats to validity, Section 6 reviews related work, Section 7 discusses key findings and implications, and Section 8 concludes the study.
@@ 2 Background
In this section, we introduce the fundamental concepts essential for understanding our paper. We cover LLMs for unit test generation along with prompt engineering, which constitutes the central ideas of this study.
2.1 LLM-based Unit Test Generation
Recent research has explored LLMs for automated test generation. Unlike traditional generators, LLMs leverage pre-trained knowledge to generate tests from natural language or code contexts. Schafer et al. (Schäfer et al., 2023) evaluated LLMs’ ability to produce valid test code. Siddiq et al. (Siddiq et al., 2024a) assessed compilability and coverage of ChatGPT-generated tests, revealing execution reliability gaps. Tang et al. (Tang et al., 2024) showed LLMs lag behind SBST tools in fault detection despite producing syntactically valid tests. Studies by Yuan et al. (Yuan et al., 2024) and Chen et al. (Chen et al., 2024) demonstrated prompt structure affects test quality. Yang et al. (Yang et al., 2024) provided cross-LLM evaluation, while Bhatia et al. (Bhatia et al., 2024) found LLM-generated Python tests suffer from incorrect assertions. Enhancement efforts include TestArt (Gu et al., 2024) and CITYWALK (Zhang et al., 2025).
Current works often focus narrowly on static correctness or coverage, lacking assessment of maintainability, runtime behavior, or test smells. Our study addresses these limitations with a holistic framework combining execution-based validation, mutation testing, and maintainability analysis. We compare five prompt engineering techniques, demonstrating how reasoning-driven prompting improves reliability, structure, and usability across three diverse datasets.
@@ 2.2 Prompt Engineering
Zero-shot learning enables models to generalize without examples but struggles with complexity (Sahoo et al., 2024). Few-shot learning uses examples to improve performance on novel tasks (Brown et al., 2020). Chain-of-Thought (CoT) (Wei et al., 2022) outlines intermediate reasoning steps, while Self-consistency (Wang et al., 2022) uses diverse reasoning paths to select consistent answers. Tree-of-Thoughts (ToT) (Long, 2023) builds on CoT by exploring a reasoning ”tree” with search algorithms. This study applies these techniques to adapt LLMs for unit test generation.
@@ 2.5 Test Smells
Unit testing verifies individual components’ functionality while maintaining readability. Test smells indicate poor practices (Palomba et al., 2016a; Peruma et al., 2019) that lead to fragility and tight coupling. Assertion Roulette and Magic Number Tests—documented in TsDetect [^1] https://testsmells.org/pages/testsmellexamples.html—commonly appear in both EvoSuite and LLM-generated tests. In Listing 1, Assertion Roulette shows multiple assertThat() calls without clarifications, obscuring failure sources.
@@ 3.1 Overview
Our methodology (Figure 2) evaluates LLM-generated test suites using ZSL, FSL, CoT, ToT, and GToT prompting techniques. To assess prompt engineering and reliability (RQ1, RQ2), we analyze structural correctness and usability using MSR, CSR, syntax correctness, and compilability. For readability and maintainability (RQ3, RQ4), we employ automated metrics including a deep learning model (Scalabrino et al., 2018) that correlates with human assessments (Sergeyuk et al., 2024), alongside complexity metrics and coding standard adherence. For code coverage (RQ5), we compare LLM tests with EvoSuite across line, instruction, and method coverage. We analyze execution outcomes (RQ6) by categorizing tests as Passed, Failed, or Timeout, using mutation testing to assess fault detection. Finally, we evaluate test smell prevalence (RQ7) to determine if reasoning-based prompting improves maintainability.
@@ 3.2 Research Questions
RQ2: How does prompt engineering influence the generation of syntactically correct and compilable test suites by LLMs? We assess how prompt engineering helps LLMs generate test suites that: ❶ adhere to syntax rules and ❷ successfully compile, enabling execution without ensuring functionality.
RQ3: Beyond syntactic correctness, how well do LLMs generate human-understandable and maintainable test suites? We evaluate if LLM-generated tests: ❶ convey test logic clearly; ❷ remain adaptable over time; and ❸ integrate seamlessly into workflows, determining if they meet practical development standards.
RQ4: How do LLM-generated test suites compare to EvoSuite and align with developer expectations in terms of readability? We investigate whether LLM tests meet developer expectations by: ❶ enhancing navigation through clear naming and structure; ❷ reducing ambiguity with well-organized presentation; and ❸ supporting collaboration. We analyze how prompting techniques influence readability using an automated model correlated with human evaluation.
@@
RQ6: How do the execution outcomes of LLM-generated test suites reflect their ability to detect faults and their applicability in real-world scenarios? We analyze test execution outcomes—(1) Passed, (2) Failed, and (3) Timeout—and use mutation testing to assess fault detection capability (Fraser and Arcuri, 2011), uncovering patterns and limitations in LLM-generated tests.
RQ7: How do LLM-generated test suites compare to EvoSuite-generated test suites in terms of test smell prevalence? We compare the absence of test smells in LLM-generated tests against EvoSuite to determine whether prompt engineering produces more maintainable tests with fewer design issues than traditional SBST techniques.

### v4

@@ Abstract
This study presents the first large-scale empirical evaluation of LLM-generated unit tests at the full class level, systematically analyzing four state-of-the-art models (GPT-3.5, GPT-4, Mistral 7B, and Mixtral 8x7B) against EvoSuite across 216,300 generated test cases targeting Defects4J, SF110, and CMD (a dataset mitigating LLM training data leakage).
We evaluate five prompting techniques—Zero-Shot Learning (ZSL), Few-Shot Learning (FSL), Chain-of-Thought (CoT), Tree-of-Thought (ToT), and Guided Tree-of-Thought (GToT)—assessing syntactic correctness, compilability, hallucination-driven failures, readability, code coverage metrics, and test smells.
Reasoning-based prompting particularly GToT significantly enhances test reliability, compilability, and structural adherence in general-purpose LLMs. However, hallucination-driven failures remain a persistent challenge, manifesting as non-existent symbol references, incorrect API calls, and fabricated dependencies, resulting in high compilation failure rates (up to 86%). Moreover, test smell analysis reveals that while LLM-generated tests are generally more readable than those produced by traditional tools, they still suffer from recurring design issues such as Magic Number Tests and Assertion Roulette, which hinder maintainability. Overall, our findings indicate that LLMs can serve as effective assistive tools for generating readable and maintainable test suites, but hybrid approaches that combine LLM-based generation with automated validation and search-based refinement are required to achieve reliable and production-ready test generation.
@@ 1 Introduction
Large Language Models (LLMs)(Brown et al., 2020; Fan et al., 2023) have emerged as a promising alternative for test generation, potentially addressing traditional approaches’ fundamental limitations. Initial studies suggest LLM-generated tests feature more natural naming conventions, human-like assertions, and better readability than SBST alternatives(Siddiq et al., 2024a; Yuan et al., 2023; Tang et al., 2024). Moreover, LLMs can generate both functional and exploratory tests beyond mere structural coverage (Wang et al., 2024a), aligning with comprehensive testing practices. Yet, it remains unclear whether prompting strategies that succeed in general code generation also transfer effectively to test generation. While reasoning-based prompts (e.g., CoT, ToT) have shown strong results in functional code tasks (Li et al., 2025; Wang et al., 2024b), unit test generation involves additional challenges: importing the class under test (CUT), setting up fixtures, managing program state, and writing reliable oracles (Molina et al., 2025; Hossain and Dwyer, 2024). These aspects introduce domain-specific constraints absent in standard code synthesis.
In this work, we examine whether prompting hierarchies generalize to test generation, where structure, dependencies, and correctness criteria differ from conventional code synthesis. Our class-level evaluation across diverse benchmarks uncovers key limitations that must be addressed before LLM-based testing can be reliably deployed:
@@
  • Structural Correctness & Extractability of LLM-Generated Tests:
    Prior work has evaluated various aspects of LLM-generated tests, including compilability and coverage (Tang et al., 2024; Yang et al., 2024; Siddiq et al., 2024a), and in some cases, test smells (Siddiq et al., 2024a), providing valuable insights into the runtime viability and maintainability of generated tests. However, few studies assess whether LLM-generated outputs are structurally valid and directly usable as standard JUnit test classes. We build on this line of work by applying extractability metrics (MSR and CSR (Macedo et al., 2024)) to measure whether generated files conform to JUnit expectations and can be seamlessly integrated into real-world testing pipelines.
@@
This paper: We systematically evaluate LLM-generated test suites and prompt engineering impact. Building on prior comparisons (Tang et al., 2024; Siddiq et al., 2024a; Yang et al., 2024), surveys (Wang et al., 2024a), and AI risk studies (Sallou et al., 2023), we conduct a large-scale evaluation across four LLMs and five prompting techniques against EvoSuite. Our study focuses on understanding prompting strategies in LLM-based test generation. We contrast LLM outputs with EvoSuite (Fraser and Arcuri, 2011), the state-of-the-art in search-based test generation (SBST) for Java, which has consistently outperformed random testing and symbolic execution (Shamshiri et al., 2015). In the SBFT Tool Competition (Jahangirova and Terragni, 2023), EvoSuite consistently ranks first in line/branch coverage and mutation score, outperforming Randoop (Pacheco and Ernst, 2007), Kex (Abdullin and Itsykson, 2022), and UTBot [^1] https://www.utbot.org/, and remains the most widely adopted and validated SBST tool (Ouédraogo et al., 2025; Almasi et al., 2017; Shamshiri et al., 2015). By positioning LLMs against EvoSuite, we ensure a strong, credible baseline (the de facto reference for automated Java test generation), providing a fair reference point for assessing whether LLMs, with proper prompting, can reach or surpass state-of-the-art SBST methods. Unlike method-level approaches in prior work, our class-level evaluation aligns with real-world testing workflows.
@@
The remainder of this paper is structured as follows: Section 2 covers LLM-based test generation, Section 3 outlines our experimental setup, Section 4 presents evaluation results, Section 5 discusses threats to validity, limitations and future work, Section 6 reviews related work, Section 7 discusses key findings and implications, and Section 8 concludes the study.
@@ 2 Background
In this section, we introduce the core concepts underpinning our study. We outline key ideas from prompt engineering, describe metrics adapted or extended for evaluating generated tests (e.g., MSR and CSR), clarify the distinction between readability and understandability, present common test smells, and review static analysis tools used to assess code quality. These notions provide essential context for interpreting the influence of prompting strategies on LLM-based test generation.
@@ 2.1 Prompt Engineering
Zero-shot learning enables models to generalize without examples but struggles with complexity (Sahoo et al., 2024). Few-shot learning uses examples to improve performance on novel tasks (Brown et al., 2020). Chain-of-Thought (CoT) (Wei et al., 2022) outlines intermediate reasoning steps, while Self-consistency (Wang et al., 2022) uses diverse reasoning paths to select consistent answers. Tree-of-Thoughts (ToT) (Long, 2023) builds on CoT by exploring a reasoning “tree” with search algorithms. This study applies these techniques to adapt LLMs for unit test generation.
@@ 2.2 Match Success Rate (MSR) and Code Extraction Success Rate (CSR)
In our setting, we adopt these metrics as a preliminary lens: rather than asking whether the code is syntactically valid or compilable, we first ask whether test code can be consistently detected and cleanly extracted from raw LLM outputs. This allows us to decouple format and extractability from later concerns of syntax, compilation, and execution.
Rationale.
LLM outputs often interleave code with natural language (e.g., explanations, caveats, reformatted snippets). This interleaving can bias downstream compilation- or execution-based metrics if extraction is not controlled. Following Macedo et al. (2024), we decouple format/extractability from syntax/compilation using two format-oriented metrics tailored to JUnit tests.
Definitions (adapted to test generation).
Let $\mathcal{O}$ be the set of all LLM responses for a given setting (model $\times$ prompt $\times$ class). Each response $o\in\mathcal{O}$ is treated as raw text. We compute:
  • MSR (Match Success Rate). Fraction of responses where a JUnit-oriented detector finds a code segment that appears to delimit a test suite. Formally,
    $\mathrm{MSR}=\frac{|\{\,o\in\mathcal{O}\mid\mathrm{detect}(o)=\mathrm{true}\,\}|}{|\mathcal{O}|}.$ (1)
    Detection succeeds when either (i) explicit delimiters are respected (“###Test START###” … “###Test END###”); or (ii) in their absence, fallback patterns match a plausible JUnit test file (package/imports/class declaration with @Test or JUnit imports). MSR is a format-compliance signal: it tells us whether a response contains something that looks like a test suite, not whether that suite is valid or compilable.
  • CSR (Code Extraction Success Rate). Fraction of responses where extraction yields a structured test file that passes minimal structural sanity checks. Formally,
    $\mathrm{CSR}=\frac{|\{\,o\in\mathcal{O}\mid\mathrm{extract}(o)\rightarrow\widehat{t}\ \wedge\ \mathrm{is\_structured}(\widehat{t})\,\}|}{|\mathcal{O}|}.$ (2)
    The extraction function isolates the best candidate $\widehat{t}$ (single file) after stripping any non-code content. $\mathrm{is\_structured}(\cdot)$ verifies that the candidate looks like a complete JUnit test file (e.g., optional package, one or more imports including JUnit, a public class with at least one method or assertion). CSR is independent of syntax or compilation (both are assessed later in RQ2).
Detectors and patterns.
We use a two-stage detector:
  1. 1.
    Delimiter-first. If the prompt requested explicit boundaries (“###Test START###” / “###Test END###”), we extract the enclosed block.
  2. 2.
    Fallback patterns. If delimiters are absent or malformed, we search for a single best candidate using ordered patterns that privilege JUnit-shaped files (package line, imports, class declaration, and at least one @Test or assertion).
If multiple code fences exist, we select the longest candidate that satisfies the JUnit shape. If no candidate passes detection, both MSR and CSR count that response as a failure.
Structural validation for CSR.
[…]
Illustrative example (Divider).
Figure 1 illustrates a typical case: the user asks for a JUnit suite for a simple Divider class. ChatGPT returns (i) a valid JUnit test class DividerTestSuite, and (ii) a natural-language summary of the scenarios. In our framework, MSR=1 because the detector identifies a plausible test file, and CSR=1 because extraction isolates a structured class with imports, assertions, and balanced braces. The trailing explanation is ignored at this stage. This example shows why MSR/CSR are necessary: without controlled detection and extraction, downstream compilation (RQ2) would be contaminated by the extraneous prose.
What MSR/CSR do not measure.
Neither metric judges Java syntax correctness, compilation, coverage, or fault detection. They are pre-execution format/extractability filters. As shown by Macedo et al. (2024), controlling output format can substantially affect downstream compilation; we therefore isolate this step as RQ1.
Synthetic corner cases.
  1. 1.
    MSR = 1, CSR = 1. Response has delimiters and encloses a full JUnit file.
  2. 2.
    MSR = 1, CSR = 0. Detector matches a region (e.g., a Markdown fence), but extracted content fails structural checks.
  3. 3.
    MSR = 0, CSR = 0. Response is purely explanatory or pseudo-code, no code region is detected.
Scope note.
MSR/CSR are LLM-specific. SBST tools (e.g., EvoSuite) directly emit JUnit files by construction; applying MSR/CSR would always succeed and thus be uninformative. We therefore report MSR/CSR only for LLM outputs and use syntax/compilation checks for both paradigms in RQ2.
@@ 2.4 Test Smells
Unit testing verifies individual components’ functionality while maintaining readability. Test smells indicate poor practices (Palomba et al., 2016; Peruma et al., 2019) that lead to fragility and tight coupling. Assertion Roulette and Magic Number Tests—documented in TsDetect [^2] https://testsmells.org/pages/testsmellexamples.html—commonly appear in both EvoSuite and LLM-generated tests. In Listing 1, Assertion Roulette shows multiple assertThat() calls without clarifications, obscuring failure sources.
@@ 4 Repository repo = executeAndWaitFor(cloneOp);
2.5 Static Analysis Tools: Checkstyle, PMD, and SpotBugs
To assess the structural quality of LLM-generated test suites, we rely on three widely used static analysis tools: Checkstyle [^3] https://checkstyle.sourceforge.io, PMD [^4] https://pmd.github.io/, and SpotBugs [^5] https://spotbugs.github.io/.
Checkstyle.
Checkstyle is a static code analysis tool that verifies Java code conformance to coding standards. It detects formatting issues, naming convention violations, and structural inconsistencies. In our study, we adopt both the Sun Code Conventions [^6] https://www.oracle.com/java/technologies/javase/codeconventions-contents.html and Google Java Style [^7] https://google.github.io/styleguide/javaguide.html profiles to evaluate coding standard adherence.
PMD.
PMD is a source code analyzer that detects common code issues such as unused variables, empty catch blocks, and overly complex code. It provides metrics for both cyclomatic complexity (control flow paths) and cognitive complexity (human-perceived difficulty). Following SonarSource guidelines [^8] https://www.sonarsource.com/docs/CognitiveComplexity.pdf, we classify cognitive complexity into: Low (<5<5), Moderate (551010), High (11112020), and Very High ($\geq 21$). For cyclomatic complexity (based on PMD thresholds), we use: Low (1144), Moderate (5577), High (881010), and Very High ($\geq 11$). These thresholds allow us to assess whether generated test cases remain structurally simple and maintainable.
SpotBugs.
SpotBugs detects potential software bugs through bytecode-level analysis. It categorizes issues based on severity into four ranked priority levels:
  • Scariest: Critical defects that threaten correctness, security, or stability.
  • Scary: Severe bugs that may cause subtle but impactful failures.
  • Troubling: Moderate issues, potentially leading to runtime problems if left unaddressed.
  • Of Concern: Minor issues or code smells that may reduce code quality but rarely break functionality.
This categorization allows us to distinguish not only the number of detected issues, but also their potential severity and impact on code reliability. By using these severity levels, we can better characterize the trade-offs between code structure and safety in automatically generated test suites.
@@ 3.1 Overview
Our methodology (Figure 2) evaluates LLM-generated test suites using ZSL, FSL, CoT, ToT, and GToT prompting techniques. To assess prompt engineering and reliability (RQ1, RQ2), we analyze structural correctness and usability using MSR, CSR, syntax correctness, and compilability. For readability and maintainability (RQ3, RQ4), we employ automated metrics including a deep learning model (Scalabrino et al., 2018) that correlates with human assessments (Sergeyuk et al., 2024), alongside complexity metrics and coding standard adherence. For code coverage (RQ5), we compare LLM tests with EvoSuite across line, instruction, and method coverage. Finally, we evaluate test smell prevalence (RQ7) to determine if reasoning-based prompting improves maintainability.
@@ 3.2 Research Questions
RQ0: What is the ability of SBST (EvoSuite) and LLMs to generate test code at all? Before analyzing correctness, readability, or maintainability, it is essential to first assess whether different paradigms can actually produce test code for a given dataset. We therefore examine the raw generation capacity of EvoSuite and LLMs across datasets, measuring how many test suites were produced out of the total expected runs. This provides a baseline for understanding the feasibility of test generation before deeper analyses.
@@
RQ2: How does prompt engineering influence the generation of syntactically correct and compilable test suites by LLMs, in comparison with SBST (EvoSuite)? We assess how different prompting strategies affect the ability of LLMs to generate test suites that ❶ adhere to syntax rules and ❷ successfully compile, enabling execution without ensuring functionality. In addition, we contrast these results with EvoSuite as a representative SBST baseline, in order to highlight similarities and differences between search-based and LLM-based test generation with respect to syntactic validity and compilability.
RQ3: Beyond syntactic correctness, how well do LLMs generate human-understandable and maintainable test suites? We evaluate if LLM-generated tests: ❶ convey test logic clearly; ❷ remain adaptable over time; and ❸ integrate seamlessly into workflows, determining if they meet practical development standards. EvoSuite is not included in RQ3 analyses, as its primary goal is maximizing structural coverage rather than producing understandable or maintainable tests. This distinction has been repeatedly emphasized in the literature: Rojas et al. (2015) show that while EvoSuite increases coverage and reduces testing time, developers consistently struggle with the readability of its outputs. Similarly, Shamshiri et al. (2018) report that although developers are as effective with EvoSuite-generated tests as with manual ones, they take significantly longer to maintain them and perceive them as harder to understand. These findings confirm that readability and understandability are not baselines EvoSuite provides, but rather limitations that have motivated recent efforts to refactor generated tests with LLMs (Deljouyi et al., 2025; Biagiola et al., 2025). Therefore, RQ3 focuses exclusively on LLM-generated tests, where prompt design explicitly targets human-oriented quality dimensions such as readability and maintainability.
RQ4: How do LLM-generated test suites compare to EvoSuite and align with developer expectations in terms of readability? We investigate whether LLM tests meet developer expectations by: ❶ enhancing navigation through clear naming and structure; ❷ reducing ambiguity with well-organized presentation; and ❸ supporting collaboration. We analyze how prompting techniques influence readability using an automated model correlated with human evaluation. While EvoSuite was excluded in RQ3 because static analysis tools such as Checkstyle, PMD, and SpotBugs assess dimensions (style compliance, cyclomatic/cognitive complexity, and bug patterns) that it was never designed to optimize, making a direct comparison unfair and potentially misleading, in RQ4 we include EvoSuite as a baseline reference. Here, the focus is on readability from a developer perspective, evaluated with an established readability model (Scalabrino et al., 2018). EvoSuite’s widespread adoption in SBST makes it a useful empirical point of comparison to determine whether LLM-generated tests deliver tangible gains in human-oriented quality, complementing the structural strengths of SBST with improved clarity and readability.
@@
RQ6: How do LLM-generated test suites compare to EvoSuite-generated test suites in terms of test smell prevalence? We compare the absence of test smells in LLM-generated tests against EvoSuite to determine whether prompt engineering produces more maintainable tests with fewer design issues than traditional SBST techniques.